Solve for the determinant in the equation below. (10)

 

 1.7.1. 

4 −3 2

1 2 −2

2 −1 −4

 1.7.2

2 −2 1

2 2 1

4 1 3


1
Expert's answer
2022-05-20T07:18:00-0400

1.7.1.

Let's solve for the determinant:


"\begin{vmatrix}
 4 & -3 & 2 \\
 1 & 2 & -2 \\
 2 & -1 & -4
\end{vmatrix}" = 4*2*(-4) + 1*(-1)*2 + (-3)*(-2)*2 - (2*2*2 + (-3)*1*(-4) + (-2)*(-1)*4) = -50


Answer: the determinant is equal to -50.


1.7.2.

Let's solve for the determinant:


"\begin{vmatrix}
 2 & -2 & 1 \\
 2 & 2 & 1 \\
 4 & 1 & 3
\end{vmatrix}" = 2*2*3 + 4*(-2)*1 + 2*1*1 - (4*2*1 + 2*(-2)*3 + 2*1*1) =8


Answer: the determinant is equal to 8.